OpenStudy (anonymous):

2cos^2(x)-3cos(x)=-1 solve for x in[0,2pi) that is 0 lessen or eual to x <2pi

OpenStudy (campbell_st):

this is an equation that can be reduced to a quadratic if you rewrite the equation you get \[2\cos^2(x) - 3\cos(x) + 1 = 0\] which is an equation that can be factored..